In order to buy a car, you borrow $33,000 from a friend at 8%/year compounded monthly for 4 years. You plan to repay the loan with 48 equal monthly payments.

What is the remaining balance after the 37th payment?
Answer #1

Formula to calculate EMI with compound interest = [P * R * (1+R)N ] / [(1+R)N -1]

where P = $33,000

R = 8% per year which would be 8% / 12 = 0.0066 monthly

N = 48 months

EMI = [33,000 * 0.0066 * (1+0.0066)48] / [(1+0.0066)48 - 1]

= 298.66 / 0.3712 = 804.57

After 37th payment = 37 * 804.57 = 29,769.45 is repaid.

With 48 installments, you are liable to pay 804.57 * 48 = 38,619.36 out of which 29,769.45 is repaid. 38,619.36 - 29,769.45 = 8,849.91 is pending after 37th installment.